# I2VControl: Disentangled and Unified Video Motion Synthesis Control

### v2

@@ Abstract
Video synthesis techniques are undergoing rapid progress, with controllability being a significant aspect of practical usability for end-users. Although text condition is an effective way to guide video synthesis, capturing the correct joint distribution between text descriptions and video motion remains a substantial challenge. In this paper, we present a disentangled and unified framework, namely I2VControl, that unifies multiple motion control tasks in image-to-video synthesis. Our approach partitions the video into individual motion units and represents each unit with disentangled control signals, which allows for various control types to be flexibly combined within our single system. Furthermore, our methodology seamlessly integrates as a plug-in for pre-trained models and remains agnostic to specific model architectures. We conduct extensive experiments, achieving excellent performance on various control tasks, and our method further facilitates user-driven creative combinations, enhancing innovation and creativity. The project page is: https://wanquanf.github.io/I2VControl. $\dagger$ Corresponding author
@@
Figure 1: We propose I2VControl, an all-in-one unified framework for image-to-video motion synthesis control. In the illustration, we show several scenarios of disentangled controls, including camera movement (camera dollies in and gets closer to the sculpture), object movement (the astronaut walks forward) and motion brush (smoke flows in the wind, with a given motion strength value). Users can select the control modes according to their requirements, where the control modes can be combined without conflict.
@@ 1 Introduction
Video synthesis technology has significant potential across various fields, such as film production, education, gaming, social media, and advertising. Recently, video synthesis technology has made rapid advancements, with state-of-the-art models like Sora [5], Kling [21] and Vidu [1] achieving outstanding results in generating high-quality video content. In this context, controllability has become crucial for optimizing video production workflows and enhancing content quality, allowing creators to precisely adjust and present their desired visual narratives.
@@
While the integration of text and visual modalities [27] is widely adopted, relying solely on textual conditions to guide video creation poses a significant challenge: the ability to accurately align textual descriptions with the complex motions in videos. This misalignment can lead to a synthesis that strays far from user intentions, ultimately constraining the practical applications of the technology. To address this challenge, prior approaches have attempted to guide the video synthesis process by injecting additional control signals, such as motion drag based control [38, 33] and camera pose control [36, 14]. These approaches, however, often focus on a single type of motion pattern and necessitate the construction of specific data formats tailored for certain scenarios, lacking a comprehensive framework for the integration of multiple control modalities. This limitation proves to be a significant barrier for complex video synthesis tasks. Consequently, these methods, while innovative, do not fully meet the expectations for a satisfactory and intuitive user experience. They often lack the versatility and flexibility required to seamlessly control the video synthesis process, thus providing users with tools that may not be sufficiently adaptable to the wide range of creative demands encountered in video production.
In this paper, we propose a novel video control framework, I2VControl, designed to integrate and manage multiple control signals within a single cohesive system. Our framework addresses the limitations of previous methods, which often focus on a single type of motion pattern and require specific data forms. To achieve comprehensive control, we conceptualize the video into trajectory function on individual motion units, each represented by disentangled control signals. Specifically, we categorize the motion units into three types: borderland, drag-units, and brush-units. The borderland units are nearly static, serving as a background with minimal motion. Drag-units are designed for more complex and flexible control, allowing users to manipulate them with six degrees of freedom for precise movement. Brush-units enable control through motion strength values, offering a simpler, yet effective way to influence motion. Trajectory function on each motion unit is further decomposed into a linear term and an additional term, where the decomposition strategy depends on the unit category. By leveraging this detailed structure, I2VControl ensures that each part of the video can be controlled with the desired granularity, from broad strokes to intricate adjustments, meeting the diverse needs of video creators.
For the network structure, we employ a plug-in architecture that equips the video synthesis model with the capability to fit the aforementioned user interaction process. We also devise a data pipeline to align the training videos with our video representation, streamlining the training process. Our approach allows users to specify various control modalities without logical conflicts, such as camera control, motion brush control, and object dragging, within a unified framework. This flexibility not only enhances the user experience by providing intuitive and precise control but also enables the exploration of combined control methods for richer and more creative video synthesis. Experiments show that our framework achieves excellent flexibility and effectiveness across a diverse range of control scenarios. In summary, our contributions include:
  • We propose I2VControl, a video motion control framework that unifies multiple control tasks.
  • We conceptualize video as motion units and expand trajectory functions on each unit as control signals.
  • We develop a plug-in network architecture and a complete data pipeline to conduct the training process.
  • Experiments show that our framework not only achieves excellent flexibility and effectiveness but also stimulates user creativity in video production.
@@ 2.1 Text to Video Synthesis
Text-to-video (T2V) generation is a challenging task that requires both single-modal realism and cross-modal consistency. In the early stage, researchers attempted to solve this task using variational auto-encoders [32], generative adversarial networks [22], and sequential generative models [40, 39, 18, 9]. Despite the progress they made, the performance of these methods was still far from expectations. Recently, diffusion probabilistic models (DPMs) have emerged and boosted the progress of video generation, achieving unprecedented high-quality generation results [5, 3]. Imagen Video [17] and Make-A-Video [30] first generated short videos with low resolutions from texts and then performed spatial super-resolution and temporal interpolation. To reduce the dimensionality of video data, several works [4, 46, 15, 12] introduced video auto-encoders to map videos into low-dimensional latent spaces to reduce computational complexity. VideoLDM [4] extended image auto-encoder to videos. Phenaki [31] proposed a variable-length video auto-encoder using causal attention. Given the lack of high-quality text-video datasets, other researchers conducted text-to-video generation by first generating an image from text and then synthesizing videos from images [34, 10, 23]. MagicVideo-V2 [34] integrated a text-to-image generation model with a video motion generator, reference image embedding module, and frame interpolation module to construct the T2V pipeline. For long T2V generation, StreamingT2V [16] introduced short-term and long-term memory blocks to enhance video consistency, while NUWA-XL [45] combined a global diffusion to generate coarse video and a local diffusion to generate fine video content, thus obtaining extremely long video generation.
@@ 2.2 Image to Video Synthesis
Unlike text-to-video (T2V) generation, image-to-video (I2V) generation leverages a static image to define the semantic content while using text prompts to specify and control the motion dynamics within the generated videos. Several approaches have been proposed to maintain the visual fidelity of the original image while introducing realistic and contextually appropriate motion. Most straightforwardly, VideoCrafter1 [6] integrated CLIP embeddings of the static image into DPMs through a dual cross-attention layer akin to IP-Adapter [6]. Meanwhile, I2V-Adapter [11] advanced this technique by establishing interactions between the given image and the subsequent noisy frames via a cross-frame attention mechanism. Notably, Animate Anyone [19] pointed out the loss of fine-grained detail information in CLIP embeddings and creatively adopts an identical network to the DPMs for image condition injection. In contrast, DreamVideo [37] attempted to achieve high-fidelity image-to-video generation by a ControlNet-like structure. Besides, SEINE [7], PixelDance [47], and PIA [49] have independently opted to expand the input channels of DPMs for injecting image information through concatenating the additional image latent with the initial noisy latents, which can better maintain the detail fidelity theoretically and practically. Moreover, several works [41, 48, 2] combined various image condition injection approaches to achieve more fine-grained detail control.
@@ 2.3 Video Synthesis Controllability
With the development of video generation technology, some video control technologies have also been proposed recently. For camera control, MotionCtrl [36] and CamereCtrl [14] employed an adaptive encoder for the extrinsic matrix of each frame to control the camera movement, and CamCo [42] integrated an epipolar attention module to enforce the 3D-consistent constraints. I2VControl-Camera [8] proposed to control the camera movement with dense trajectory, where a motion strength controller is employed together for adjustive motion dynamic. There are several strategies for object motion control. Boximator [33] and Direct-A-Video [43] utilized the trajectory of the bounding box of the object to guide the object motion. DragNVWA [44] employed some vector arrows to guide the object movement in the input image. DragAnything [38] further toke use of the explit segmentation infomation constructed by SAM [20] to map the dragging arrow to correct pixel rigions. MOFA-Video [25] employed a sparse-to-dense strategy to convert the user input dragging arrows into dense flow and thus control the video generation process. Some polular video generation tools, such as Gen-2 [28], and Kling [21] also involved the above functions. In this work, we propose to unify these functions in one single framework, for completely disentagled control.
@@ 3 Method
Figure 2: Data pipeline of I2VControl. A video is represented as a 3D point trajectory function within the camera system at the initial frame. We perform motion unit partitioning and define the unit-wise trajectory function. Except for the borderland, other units are randomly assigned as either drag- or brush-units. The final control signal includes point trajectory, motion strength, partition and category map.
In this section, we present details of the I2VControl framework. We introduce our video representation in Sec. 3.1, the control signal construction in Sec. 3.2, our network architecture, training and inference process in Sec. 3.3.
3.1 Video Representation
In this section, we introduce the video representation used in our paper. We first introduce the motion unit partition, and then define the unit-wise trajectory function.
Motion Units. A video can be conceptualized as a sequence of frames in time range $\lambda\in[0,T]$, each capturing a projection of a deformable underlying 3D world through a camera with potential camera movement. We denote the domain of 3D points captured by the first frame (at time moment $\lambda=0$) as $\Omega\subseteq\mathbb{R}^{3}$. Typically, the domain captured by the camera is composed of multiple independent motion parts instead of one total part, which encourages us to divide the entire domain $\Omega$ into multiple independent “motion units”:
@@ 3.1 Video Representation
where $p\in\{0,1,...,P\}$ denotes the unit index, $\Omega^{(p)}$ denotes the $p$-th unit, $\chi^{(p)}$ denotes the mask of the $p$-th unit. Based on the properties and control situations, we summarize these units into three categories, defined as: brush-units, drag-units, and the borderland. Both brush-units and drag-units belong to the foreground, and the difference between them lies in the control process. The brush-units are controlled by scalar value motion strength only, while the drag-units can be controlled by $6$-DOF ($6$ degrees of freedom) motion guidance (a rigid transformation within the group $SE(3)$) and additional motion strength. During training, we randomly set a foreground part as one of the two categories. During inference process, the choice is made by the users. The region that are not covered by the brush-units and the drag-units is summarized as a single part, which we call the borderland. During the training and inference process, we always consider the borderland part as the first part with index 00, both the brush-units and the drag-units are assigned indices greater than 00.
Unit-wise Trajectory Function. We have introduced the motion unit partition of the captured domain $\Omega$. Now, we construct the trajectory function on $\Omega$ to characterize the motion of the 3D points within it. Without loss of generality, we use the camera coordinate system of the first frame as the world coordinate system. So we can define a trajectory function in the world coordinate system, $\mathcal{D}(\lambda,\mathbf{x}):[0,T]\times\Omega\to\mathbb{R}^{3}$, where $\mathcal{D}(\lambda,\mathbf{x})$ represents the position of the point $\mathbf{x}\in\Omega$ at time $\lambda$. At the initial moment ($\lambda=0$), the function is simply $\mathbf{x}$ itself, i.e., $\mathcal{D}(0,\mathbf{x})=\mathbf{x}$.
To better handle the unit-wise motion, on each unit $\Omega^{(p)}$, we decompose the trajectory function into two terms: a rigid term and an additional term:
@@
Further, we convert the trajectory function into the camera coordinate system. At each time moment $\lambda$, we denote the camera extrinsic as $\mathcal{E}_{\lambda}\in SE(3)$. The transformation from the world coordinate system (i.e., the coordinate system of the first frame) to the camera coordinate system at time $\lambda$ is given by the inverse of $\mathcal{E}_{\lambda}$, denoted as $\mathcal{E}^{-1}_{\lambda}$. Thus, the transformation of the trajectory function $\mathcal{D}$ from the world coordinate system to the camera coordinate system can be expressed as:
where $\mathcal{F}(\lambda,\mathbf{x})$ represents the position of the point $\mathbf{x}$ in the camera coordinate system at time $\lambda$. Considering that $\mathcal{E}_{0}=\mathcal{I}$, we can deduce that:
which proves that $\mathcal{F}(\lambda,\mathbf{x})$ is still a trajectory function representing the point motions from their initial positions. So the decomposition in Eq. 2 becomes:
$\mathcal{F}(\lambda,\mathbf{x})$ $=\mathcal{E}_{\lambda}^{-1}\circ\mathcal{R}^{(p)}_{\lambda}\circ\mathbf{x}+\mathcal{E}_{\lambda}^{-1}\circ\mathcal{G}^{(p)}(\lambda,\mathbf{x}),\;\forall\mathbf{x}\in\Omega^{(p)},$ (5)
@@ 3.2 Control Signal Construction
In this section, we describe how we construct the control signals. We start by defining a general motion strength function in a general setting. Consider an arbitrary domain $\Gamma\subset\mathbb{R}^{3}$, and an arbitrary trajectory function $\mathcal{H}(\lambda,\mathbf{x}):[0,T]\times\Gamma\to\mathbb{R}^{3}$, we refer to I2VControl-Camera [8] and define the motion strength function $\mathcal{M}$ as:
[…]
where the result is a scalar value. An important property to note is that $\mathcal{M}$ is invariant with respect to rigid transformations on $\mathcal{H}$:
$\mathcal{M}(\Gamma,\mathcal{E}\circ\mathcal{H},\lambda)\equiv\mathcal{M}(\Gamma,\mathcal{H},\lambda),\;\forall\mathcal{E}\in SE(3),$ (7)
which implies that the motion strength is independent of the choice of coordinate system.
Now we describe how we set the $\mathcal{R}^{(p)}_{\lambda}$ and motion strength value $m_{\lambda}^{(p)}$ on different kinds of units:
@@
$\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I},\;m_{\lambda}\equiv 0.$ (8)
@@
$\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I},\;m_{\lambda}=\mathcal{M}(\Omega^{(p)},\mathcal{G}^{(p)},\lambda).$ (9)
Similar to the borderland, we set $\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I}$ on brush-units because we do not require users to provide drag controls for these parts. Instead, users only need to provide the value of $m_{\lambda}$. For the training data, we compute the motion strength with Eq. 6, which indicates the motion deviation from the identity transformation $\mathcal{I}$. During inference, users can designate an object by drawing a mask and make it move by simply classifying the region as a brush-unit and providing a motion strength value.
@@
$m_{\lambda}$ $=\mathcal{M}(\Omega^{(p)},\mathcal{G}^{(p)},\lambda).$ (11)
We set $\mathcal{R}^{(p)}_{\lambda}$ as the rigid transformation that best “fits” the trajectory, which can be obtained by solving a least-squares problem. In other words, $\mathcal{R}^{(p)}_{\lambda}$ serves as a satisfactory “proxy” for the motion in this part. Considering its rigid nature, which can be described using just 6 degrees of freedom, it is also user-friendly. The motion strength $m_{\lambda}$ is computed with the same formulation as the brush-units. However, for the drag-units, $m_{\lambda}$ indicates the motion deviation from the rigid fitting result instead of the identity transformation. During inference, users can control the motion of an object by drawing its mask, designating the region as a drag-unit, and providing both the rigid transformation and the additional motion strength value.
In addition to the unit-wise definitions, another important control aspect is the camera control, denoted as $\mathcal{E}_{\lambda}$. This control is shared by all units and is also represented as a 6-DOF signal, ensuring user-friendliness. Inspired by I2VControl-Camera, we do not directly employ $\mathcal{E}_{\lambda}$ and $\mathcal{R}^{(p)}_{\lambda}$ as the control signal. Instead, we use them to compute the point trajectory in the camera coordinate system:
@@
$\mathbf{M}_{\lambda}=\bigsqcup\limits_{p=0}^{P}(m_{\lambda}\cdot(\chi^{(p)}\odot\Omega)),\;\lambda\in[0,T],$ (13)
To pass the unit index into the network, we further construct a partition map as:
@@
Furthermore, it is essential for the network to recognize the category of each unit. We assign $c^{(p)}$ the values 0, 1, and 2 when the $p$-th unit is classified as borderland, drag-part, and brush-part, respectively. Thus, we define a category map as:
@@
Finally, we concatenate the aforementioned control signals together to form the ultimate control signal $(\mathbf{T}_{\lambda},\mathbf{M}_{\lambda},\mathbf{P}_{\lambda},\mathbf{C}_{\lambda})$.
In our data pipeline, we first apply SAM [20] to partition the first frame image into individual parts. Following the method for obtaining the dynamic mask from I2VControl-Camera, we then extract parts that intersect with the dynamic mask with an intersection area occupying more than half of the total area. These extracted parts are used as our training units, randomly set as drag-unit or brush-unit. All unselected parts are merged into a single borderland. We adhere to the data strategy employed in I2VControl-Camera, tracking $\mathcal{F}(\lambda,\mathbf{x})$ and solving $\{\mathcal{E}_{\lambda}\}_{\lambda\in[0,T]}$, and then compute $\mathcal{D}(\lambda,\mathbf{x})=\mathcal{E}_{\lambda}\circ\mathcal{F}(\lambda,\mathbf{x})$. After discretizing the aforementioned variables, our control signal is ultimately represented as a tensor with the shape $(T,5,H,W)$.
3.3 Network, Training and Inference
As mentioned above, we apply an adaptive strategy to design the network structure. Taking the $(T,5,H,W)$-shaped tensor as input, we first apply several convolutional layers to convert the input to tokens and then concatenated them with the tokens in the original diffusion process before self-attention computation. After computing self-attention, the additional parts added during concatenation are removed mannualy to restore the original token shape, following the approach of  [19]. The restored token is added back to the original diffusion token, thus completing the adapter calculation process. Except for inserting the above adapter structure, we keep all other training processes consistent with the original base model, including the loss and scheduler.

### v3

@@ Abstract
Motion controllability is crucial in video synthesis. However, most previous methods are limited to single control types, and combining them often results in logical conflicts. In this paper, we propose a disentangled and unified framework, namely I2VControl, to overcome the logical conflicts. We rethink camera control, object dragging, and motion brush, reformulating all tasks into a consistent representation based on point trajectories, each managed by a dedicated formulation. Accordingly, we propose a spatial partitioning strategy, where each unit is assigned to a concomitant control category, enabling diverse control types to be dynamically orchestrated within a single synthesis pipeline without conflicts. Furthermore, we design an adapter structure that functions as a plug-in for pre-trained models and is agnostic to specific model architectures. We conduct extensive experiments, achieving excellent performance on various control tasks, and our method further facilitates user-driven creative combinations, enhancing innovation and creativity. Project page: https://wanquanf.github.io/I2VControl.
@@
Figure 1: We propose an all-in-one disentangled and unified framework for image-to-video motion synthesis control, named I2VControl. In the illustration, we show several scenarios of controls, including camera movement (camera dollies in and gets closer to the sculpture), motion brush (smoke flows in the wind, with a given motion strength value), object movement (the astronaut walks forward), and the combination of all above control types (camera tilts down; drag the apple; brush the squirrel). Users can select the control modes according to their requirements, where the control modes can be combined without conflict.
@@ 1 Introduction
Motion controllability plays a pivotal role in video synthesis technology [5, 22, 1, 32], particularly when video creators require meticulously designed spatiotemporal dynamics. To align generated video motion with human intent — a goal often unmet by text prompts [28] due to their limited expressiveness — some approaches have attempted to guide the video synthesis process by injecting additional control signals, such as motion dragging [40, 35, 20], camera pose [38, 15] and motion brush [30, 29].
Previous motion control approaches often focus on a single type of motion pattern and necessitate the construction of specific data formats (e.g. 3D static dataset RealEstate10K [51] for camera control) tailored for certain scenarios, lacking a comprehensive framework for the integration of multiple control types. Simply integrating them into a single inference pipeline would lead to logical conflicts. For instance in Fig. 2 (a), the dragging control requires source and target positions for objects; however, if camera control is added, the target position may differ across different camera views, which could lead to inconsistencies of user intent. Another example in Fig. 2 (b) involves using a motion brush (animate a selected area while keeping other areas fixed); if camera control is added, background pixels would move, creating logical conflict.
@@
This limitation is a significant barrier for complex video synthesis tasks. Consequently, previous single-typed methods, while innovative, do not fully meet the expectations for a satisfactory and intuitive user experience. They often lack the versatility and flexibility required to seamlessly control the video synthesis process, thus providing users with tools that may not be sufficiently adaptable to the wide range of creative demands encountered in video production.
In this paper, we propose a novel video control framework, I2VControl, designed to integrate multiple control signals within a single cohesive system. To achieve this, we progressively implement three important designs:
$\bullet$ Consistent Representation. Camera control methods typically use the extrinsic matrix [38, 15] as signal, while dragging utilize sparse points [40] or bounding boxes [35, 20] to dictate object movement, and motion brush task often employs optical flow [30]. These representations are entirely disparate, making simultaneous control extremely challenging. Thus, our first step is to unify the representation across these tasks. Specifically, we choose dense point trajectories as our unified representation. In Sec. 4, we will show how we design different formulations of this representation to manage different control types. This unification allows us to integrate different tasks more easily.
$\bullet$ Spatial Partitioning. A natural observation is that each control intent correlates with a specific spatial region on the input image. The motion brush is applied to a designated area, the object being dragged has its own mask, and camera movement involves a mask formed by all areas not occupied by other controls. During inference, it is easy for users to select these masks, either manually or by using SAM [21]. With this mechanism, the framework can support multiple regional controls on one single input image.
$\bullet$ Network Structure. Despite a consistent representation and partitioning mechanism, a key challenge remains: inputting this information to the network. We need to provide the model with some key information: the partition map, the type for each region, and the control signals for each region. Thanks to the unified representation, each pixel is associated with its point trajectory, which is a regular form that can be easily inputted into the network. Additionally, we convert the partition and control types of each unit into spatial maps, which are concatenated with the trajectory to injected into the network. After several layers of convolution, the input is encoded into tokens whose dimensions is harmonious with the pretrained model. Then we employ some additional trainable attention layers to form an adapter.
Certainly, we would face a significant challenge in effectively simulating user inputs in the training data. Therefore, we devise a data pipeline to align the training videos with our designed control signal, streamlining the training process. Experimental results demonstrate that our framework offers outstanding flexibility and effectiveness across a range of control scenarios. Moreover, it matches or surpasses the top performances of previous methods in one-typed tasks, including dragging, camera control, and motion brush. In summary, our contributions include:
  • As far as we know, I2VControl is the first framework that can conduct camera control, object dragging and motion brush together in one single pipeline without conflicts.
  • To achieve unified control, we have developed the consistent representation, spatial partitioning mechanism, the adapter network, and a data pipeline for training.
  • Quantitative and qualitative results are excellent.
Figure 2: Examples of control conflicts in multiple control tasks. In (a), we try to drag the cat and apply camera tilt-up at the same time. We drag the cat towards the “right + down” direction; however, in the tilt-up view, the target position become much higher than the cat, which is an obfuscation. In (b), we try to employ motion brush and camera tilt-down together. We brush the human region, keeping the background fixed; however, in the tilt-down view, every background pixel moves, which is also a conflict.
@@ 2.1 Text to Video Synthesis
Early research on text-to-video (T2V) generation utilized VAE [34], GAN [23], and sequential generative models [42, 41, 18, 10] with limited quality. Recent advancements have been driven by diffusion probabilistic models (DPMs), significantly enhancing video quality [5, 3]. Initiatives like Imagen Video [17] and Make-A-Video [31] achieved initial low-resolution video outputs from text, which were refined through spatial super-resolution and temporal interpolation. To manage computational complexity, video auto-encoders [4, 47, 16, 13] encoded videos into compact latent spaces. Notably, Phenaki [33] developed a variable-length video auto-encoder utilizing causal attention. A hybrid approach [36, 11, 24] combined text-to-image and image-to-video synthesis to improve the video quality.
@@ 2.2 Image to Video Synthesis
Image-to-video (I2V) generation converts static images into videos. VideoCrafter1 [6] used a dual cross-attention layer to integrate CLIP embeddings of images in DPM. I2V-Adapter [12] built on this approach by facilitating interactions between the initial image and noisy frames via cross-frame attention. Animate Anyone [19] addressed the loss of detail in CLIP embeddings by using a network identical to DPMs for image condition injection. DreamVideo [39] used a ControlNet-like structure for high-fidelity outcomes. Additionally, SEINE [7], PixelDance [48], and PIA [50] expanded DPM input channels by combining image latents with noisy latents to preserve details. DynamicRafter [44], I2VGen-XL [49], and SVD [2] mixed various image condition injection techniques to handle precise details.
@@ 2.3 Video Synthesis Controllability
With the development of video generation, some motion control technologies have also been proposed recently. For camera control, MotionCtrl [38] and CamereCtrl [15] employed an adapter for the extrinsic matrix of each frame to control the camera movement, and I2VControl-Camera [8] proposed a motion strength controller for further adjustive motion dynamic. Boximator [35] and Direct-A-Video [45] utilized the trajectory of the bounding box of the object to guide the object motion. DragNVWA [46], DragAnything [40], MOFA-Video [26] utilized sparse control points to guide object motion. Motion brush functionality enables users to drive motion in specific regions via a provided mask, with a scalar motion strength value, which is supported in works like Motion-I2V [30] and Gen-2 [29]. We propose to unify multiple types in one single framework.
3 Preliminary and Rethinking
3.1 Dual Task
We first analyze two key observations regarding controllable generation, which is inspiring for algorithm design:
$\bullet$ Every controllable generation task has a perception task as its dual task. We show examples in Fig. 3, including “text-controlled generation and captioning,” “point dragging and point tracking,” and “box control and box detection.” Other instances encompass “camera pose control and 6-DoF visual odometry,” as well as “motion brush control and motion strength analysis,” among others.
$\bullet$ Dual perception task serves as data pipeline for controllable generation task. For example, we can apply captioning to obtain text-video pairs to train T2V, or track points in training videos to construct training data for dragging framework. For other tasks, the logic is similar. This observation inspires us to approach from a perception perspective, starting with a video to design a pipeline that exactly serves as the dual task to our unified control.
Figure 3: Examples of dual tasks, where the dual perception task serves as the data pipeline of the controlling generation task.
3.2 Dense Trajectory Representation
As described in Sec. 1, we choose dense trajectory as our control representation. We conceptualize video and model trajectory function inspired by I2VControl-Camera [8]. A video can be conceptualized as a sequence of frames in time range $\lambda\in[0,T]$, each capturing a projection of a deformable 3D world through a camera with potential camera movement. We denote the domain of 3D points captured by the first frame (at time moment $\lambda=0$) as $\Omega\subseteq\mathbb{R}^{3}$.
$\bullet$ Trajectory Function in World System. We use the camera coordinate system of the first frame as the world coordinate system. So we can define a trajectory function in the world coordinate system, $\mathcal{D}(\lambda,\mathbf{x}):[0,T]\times\Omega\to\mathbb{R}^{3}$, where $\mathcal{D}(\lambda,\mathbf{x})$ represents the position of the point $\mathbf{x}\in\Omega$ at time $\lambda$. At the initial moment ($\lambda=0$), we have $\mathcal{D}(0,\mathbf{x})=\mathbf{x}$.
$\bullet$ Trajectory Function in Camera System. Further, we convert the trajectory function into the camera coordinate system. At each time $\lambda$, we denote the camera extrinsic as $\mathcal{E}_{\lambda}\in SE(3)$. The transformation from the world coordinate system (i.e., the coordinate system of the first frame) to the camera coordinate system at time $\lambda$ is given by the inverse of $\mathcal{E}_{\lambda}$, denoted as $\mathcal{E}^{-1}_{\lambda}$. Thus, the transformation of the trajectory function $\mathcal{D}$ from the world coordinate system to the camera coordinate system can be expressed as:
where $\mathcal{F}(\lambda,\mathbf{x})$ represents the position of $\mathbf{x}$ in camera system at time $\lambda$. Considering $\mathcal{E}_{0}=\mathcal{I}$, we can deduce that:
which proves that $\mathcal{F}(\lambda,\mathbf{x})$ is still a trajectory function representing the point motions from their initial positions.
$\bullet$ Motion Strength. We refer to I2VControl-Camera [8] and define a general motion strength function. Consider an arbitrary domain $\Gamma\subset\mathbb{R}^{3}$ and trajectory function $\mathcal{H}(\lambda,\mathbf{x}):[0,T]\times\Gamma\to\mathbb{R}^{3}$, we define the motion strength $\mathcal{M}$ as:
[…]
yielding a scalar that denotes the average speed of points.
@@ 4 Method
In this section, we present our method details. We introduce our video representation and partitioning in Sec. 4.1, the control signal construction in Sec. 4.2, our dual task pair (data pipeline; controllable inference) in Sec. 4.3, our network and training process in Sec. 4.4. Note: In this section, we continue to use the notation established in Sec 3.2.
4.1 Video Representation and Partitioning
We first introduce the motion unit partitioning, and then define the unit-wise trajectory function formulation.
$\bullet$ Motion Units. Typically, the domain captured by the camera is composed of multiple independent motion parts instead of one total part, which encourages us to divide the entire domain $\Omega$ into multiple independent “motion units”:
@@ 4.1 Video Representation and Partitioning
where $p\in\{0,1,...,P\}$ denotes the unit index, $\Omega^{(p)}$ denotes the $p$-th unit, $\chi^{(p)}$ denotes the mask of the $p$-th unit. Based on the properties and control situations, we summarize these units into three categories, defined as: brush-units, drag-units, and the borderland. Both brush-units and drag-units belong to the foreground, and the difference between them lies in the control type. The brush-units are controlled by scalar value motion strength only, while the drag-units can be controlled by 666-DOF (666 degrees of freedom) motion guidance (a rigid transformation within the group $SE(3)$) and additional motion strength. During training, we randomly set a foreground part as one of the two categories. During inference, the choice is made by the users. The region not covered by the brush-units and the drag-units is summarized as a single part called borderland. During the training and inference process, we always consider the borderland part as the first part with index 0, both the brush-units and the drag-units are assigned indices greater than 0.
$\bullet$ Unit-wise Trajectory Function. To handle the unit-wise motion, on each unit $\Omega^{(p)}$, we decompose the trajectory function into two terms: a rigid term and an additional term:
@@ 4.2 Control Signal Construction
In this section, we describe how we construct the control signals. Each unit is assigned with a $\mathcal{R}^{(p)}_{\lambda}$ and motion strength $m_{\lambda}^{(p)}$. Now we introduce how we construct them:
@@
$\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I},\;m_{\lambda}^{(p)}\equiv 0.$ (6)
@@
$\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I},\;m_{\lambda}^{(p)}=\mathcal{M}(\Omega^{(p)},\mathcal{G}^{(p)},\lambda).$ (7)
Similar to the borderland, we set $\mathcal{R}^{(p)}_{\lambda}\equiv\mathcal{I}$ on brush-units. Instead, users only need to provide the value of $m_{\lambda}^{(p)}$. We compute the motion strength $\mathcal{M}$ with Eq. (3), which indicates the motion deviation from the identity transformation $\mathcal{I}$. During inference, users can designate an object by drawing a mask and make it move by simply classifying the region as a brush-unit and providing a motion strength value.
@@
$m_{\lambda}^{(p)}$ $=\mathcal{M}(\Omega^{(p)},\mathcal{G}^{(p)},\lambda).$ (9)
We set $\mathcal{R}^{(p)}_{\lambda}$ as the rigid transformation that best “fits” the trajectory, which can be obtained by solving a least-squares problem. In other words, $\mathcal{R}^{(p)}_{\lambda}$ serves as a satisfactory “proxy” for the motion in this part. Considering its rigid nature, which can be described using just 6 degrees of freedom, it is also user-friendly. The motion strength $m_{\lambda}^{(p)}$ is computed with the same formulation as the brush-units. However, for the drag-units, $m_{\lambda}^{(p)}$ indicates the motion deviation from the rigid fitting result instead of the identity transformation. During inference, users can control the motion of an object by drawing its mask, designating the region as a drag-unit, and providing both the rigid transformation and the additional motion strength value.
In addition to the unit-wise definitions, another important control is the camera control, denoted as $\mathcal{E}_{\lambda}$. This control is shared by all units and is also represented as a 6-DOF signal, ensuring user-friendliness. Finally, we can compute the point trajectory in the camera coordinate system:
@@
$\mathbf{M}_{\lambda}=\bigsqcup\limits_{p=0}^{P}(m_{\lambda}^{(p)}\cdot(\chi^{(p)}\odot\Omega)),\;\lambda\in[0,T],$ (11)
To represent the unit index, we construct a partition map:
@@
Furthermore, it is essential for the network to recognize the category of each unit. We assign $c^{(p)}$ the values 0, 1, and 2 when the $p$-th unit is classified as borderland, drag-part, and brush-part, respectively. So we define a category map:
@@
Finally, we concatenate the aforementioned tensors to form the ultimate control signal $(\mathbf{T}_{\lambda},\mathbf{M}_{\lambda},\mathbf{P}_{\lambda},\mathbf{C}_{\lambda})$.
Figure 4: Our data pipeline to convert RGB video into control signals and conduct the training. For more details, please see Sec. 4.3.
4.3 Our Dual Task Pair
As discussed in Sec. 3.1, we introduce our dual task pair.
[…]
Figure 5: The process of obtaining the motion units.
$\bullet$ Unified Control: Inference. During inference, as a user, we need to construct the control signal consistent with the output of data pipeline. We first use the Unidepth [27] metric depth estimation method to obtain the initial point set $\Omega$. Users then interact with our system to provide further input. They can first select the motion units of interest. Specifically, users can employ a bounding box and text description to select a small patch, and then set it as either a brush-unit or a drag-unit. If set as a drag-unit, users are required to provide a 6-DOF input to control the movement of that patch. Users can repeatedly select multiple patches until they have completed their design. Additionally, users can input an overall camera movement, which does not conflict with the motion of the selected patches. Our system provides a preview interface, allowing users to see the results of the drag and camera movement in real-time, achieving convenient “what you see is what you get” experience.
Figure 6: Details of adapter architecture and calculation.
4.4 Network and Training
As mentioned above, we apply an adapter network structure. We show the architecture and pseudo code of adapter in Fig. 6, where the green layers are trainable. Taking the $(T,5,H,W)$-shaped tensor as input, we first apply several convolutional layers to convert the input to tokens and then concatenated them with the tokens in the original diffusion process before self-attention computation. After computing self-attention, the additional parts added during concatenation are removed to restore the original token shape. The restored token is added back to the original diffusion token, thus completing the adapter calculation process.
While training, we extract control signal from ground truth video, then pass it as input to the adapter. Except for inserting the above adapter structure, we keep all other training settings consistent with the original base model, including the loss (only MSE loss) and scheduler.